OpenStudy (anonymous):

Evaluate the following. Write your answer in scientific notation. (3 x 107)(6.1 x 103)

OpenStudy (lgbasallote):

do you mean \[(3 \times 10^7)(6.1 \times 10^3)?\]

OpenStudy (lgbasallote):

i believe you're allowed to use a calculator right? just type in 3 (EXP) 7 x 6.1 (EXP) 3 my calculator judged it as \(1.83 \times 10^{11}\)

OpenStudy (anonymous):

 would that work?